OpenStudy (anonymous):

find all of the solutions of the following equation in the interval (0,2π). 2sinx-cscx=0

OpenStudy (anonymous):

2sinx - 1/sinx = 0 2sin^2(x) - 1 = 0 sin^2(x) = 1/2 sinx = (root2)/2 x = π/4, 3π/4